OpenStudy (koikkara):

A and B undertake to do a piece of work for Rs. 600. A alone can do it in 6 days while B alone can do it in 8 days. With the help of C, they finish it in 3 days. !find the share of each. Sol :C's 1 day's work = 1/3-(1/6+1/8)=24 A : B : C = Ratio of their 1 day's work = 1/6:1/8:1/24= 4 : 3 : 1. A’s share = Rs. (600 *4/8) = Rs.300, B's share = Rs. (600 *3/8) = Rs. 225. C's share = Rs. [600 - (300 + 225») = Rs. 75.

OpenStudy (koikkara):

How A’s share = Rs. (600 *4/8).... 4/8 came from ?

OpenStudy (koikkara):

@Compassionate hope you can help me....my brain looks sleepy...lol

OpenStudy (koikkara):

@dan815 ...thanks for your valuable time... Focus only.... A’s share = Rs. (600 *4/8) = Rs.300, B's share = Rs. (600 *3/8) = Rs. 225. C's share = Rs. [600 - (300 + 225») = Rs. 75. Where did 4/8 come from ?

OpenStudy (dan815):

A and B undertake to do a piece of work for Rs. 600. A alone can do it in 6 days while B alone can do it in 8 days. With the help of C, they finish it in 3 days. !find the share of each.

OpenStudy (dan815):

okay so A works are a rate of W/6 in 3 days he will have done half the total work as 3*w/6= w/2

OpenStudy (dan815):

in 3 days B would have done 3*w/8

OpenStudy (koikkara):

Hmm, but i don't understand how they got \(*4/8)\), in here...

OpenStudy (dan815):

it should say 3/6 but something

OpenStudy (koikkara):

lol, i check google, everywhere they got same....lol,ccrazy : http://indiah.weebly.com/chapter-15-time-and-work.html

OpenStudy (koikkara):

Ex8

OpenStudy (dan815):

just think about it like this its pretty simple once u understand

OpenStudy (dan815):

in 6 days A finishes work so in 1 day A finishes 1/6th of the work

OpenStudy (dan815):

in 8 days B finishes work so in 1 day B finishes 1/8th of the work

OpenStudy (koikkara):

Fine..

OpenStudy (dan815):

so how much of the work did these 2 finish in 3 days with themself thats 3*1/6 + 3*1/8 of the work 3/6 + 3/8 = 4/8+3/8=7/8 of the work was done by a and b so c did 1/8th of the work in 3 days

OpenStudy (koikkara):

Hey, i don't understand the money part, rest goes good.

OpenStudy (dan815):

now u split your 600 dollars up evenly between these ppl A did half the work so 300 for him B did 3/8 so 3/8*600 Cdid 1/8 sp 1/8 * 600

OpenStudy (koikkara):

thats good..

OpenStudy (dan815):

okay so u got it now?

OpenStudy (koikkara):

that was great..but still need to know 4/8 as its a time saving method for faster calculations... Thanks @dan815 !! Let me post it in facebbok...they may do that... Thanks again...lol